# Rethinking Knowledge Distillation: A Data Dependent Regulariser With a Negative Asymmetric Payoff

### v1

@@ Appendix C Feature Map Matching Knowledge Distillation
Largely we see that the results for Feature Map KD correspond to our original findings, when there is statistically significant functional transfer the transfer is asymmetric in nature and is weighted towards incorrect predictions. While there is a difference between blocks 4 and 5, understanding this fully this would require further exploration to make concrete statements about why this difference emerges.
Appendix D Random Control Distillation (RCD) Comparison to Label Smoothing
@@
We also rely on RCD, which retains soft targets but replaces the teacher’s logits with uniform noise. RCD preserves any label-smoothing benefit while removing semantic content. Across all metrics, we find that LS and RCD match or exceed KD in accuracy, yet exhibit no increase in functional similarity with the teacher, particularly on incorrect predictions. This confirms that KD’s asymmetric error transfer arises from the specific structure of the teacher’s logits, not from softening per se.
@@
Appendix E Knowledge Distillation to Smaller Student
Justification:
This setup allows for an analysis of Knowledge Distillation where the student model is smaller than the teacher model, as expected in practice.
Caveat:
Although this moves away from our traditional experiential setup where the student can perfectly match the teacher, we use this example to show how transfer works between a larger teacher to a smaller student. It is important to note that using a smaller student introduces uncertainty on if the student capacity is a bottleneck to knowledge transfer. However, given that in practice Knowledge Distillation is used in this setting we show how our fundamental insights from the self distillation case transfer to other cases of dilatation. Our study of using a smaller students is not exhaustive but demonstrative and verifies the findings presented in the main body of the paper, and the utility of our initial experimental setup. Other than the architecture’s implicit bias towards the problem, which affects its performance (loss and accuracy), there are no confounding factors that could influence Knowledge Distillation.
E.1 TinyImageNet ResNet50 Teacher to ResNet18 Student
@@
Findings:
We observe a low train loss for the teacher model circa 0.0014 with a high train accuracy circa 0.9998; see Table 15. This low train loss corresponds as expected, with no significant knowledge transfer across alpha values; see Tables 16, 17, 18 and 19. This result is as expected from the results and intuition presented in the results of the main body of the paper. It highlights how this finding generalises to the practical KD environment.
@@
| KD 0.1 | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
| KD 0.5 | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
| KD 0.9 | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
E.2 Tiny Shakespeare Nano-GPT Teacher to Pico-GPT Student
@@
The Nano-GPT Teacher is a GPT2-style transformer with an embedding dimension of 384, a vocabulary size of 65, six attention heads, six transformer blocks, a dropout of 0.200, and a block size of 256. The Pico-GPT student has an embedding dimension of 192, halving the internal width of the model; all other model settings are the same as the teacher.
The teacher and student are trained on the Tiny Shakespeare dataset, with the first 90% used for training and the last 10% used for testing. The dataset was tokenised via a character tokeniser, and the model was trained auto-regressively to predict the next character token. The teacher and student are trained with the Adam optimiser with a learning rate of 3e-4 with a batch size of 64 for 5000 iterations. The student models are trained with the same seeds and data orders from seeds 10 to 19 for the 10 models used for averaging. This is repeated for the three teachers trained on seeds 0 to 2.
Justification:
This setup allows for an analysis of Knowledge Distillation where the student model is smaller than the teacher model, as expected in practice. It is not exhaustive but demonstrative that the findings we present in the main body of the paper generalise to this case. Other than the architecture’s implicit bias towards the problem, which affects its performance (loss and accuracy), no confounding factors could influence Knowledge Distillation.
Findings:
We observe a high train loss for the teacher model circa 0.86 with a high train accuracy circa 0.72; see Table 20. This high train loss corresponds as expected with a substantial knowledge transfer which increases as alpha increases, see Tables 96, 97, 98 and 99. This substantial knowledge transfer coincides with an asymmetric payoff in prediction agreement, strongly favouring incorrect predictions, see Figure 65(a). This result is as expected from the results and intuition presented in the results of the main body of the paper and highlights how this finding generalises.

### v2

@@ 4.3 KD Variants: Feature-Map Matching Distillation
Largely, we find that the results for feature-map matching KD corroborate our original findings: when there is statistically supported functional transfer, the transfer is asymmetric in nature and weighted towards incorrect predictions. While the results differ between blocks 4 and 5, they highlight that a higher degree of functional similarity corresponds directly to increased negative asymmetric knowledge transfer.
4.4 Temperature and Negative Asymmetric Transfer
In this section we explore how temperature affects the negative asymmetric payoff of knowledge distillation. We demonstrate this on the TinyShakespeare and ImageNet datasets, which exhibit the highest levels of transfer in the vision and language domains. We consider temperature values of 1 and 2 for ImageNet, and 1, 2 and 4 for TinyShakespeare, using students that are smaller than the teachers to reflect the standard KD setup. The raw results tables for this section can be found in Appendix Section C.2.1 and C.3.1 for ImageNet and TinyShakespeare, respectively.
4.4.1 Computer Vision
Impact on Accuracy:
We find that increasing the KD temperature reduces the accuracy utility of KD, this is expected as increasing the temperature reduces the signal between the student and the teacher. In table 25 it can be observed that using a temperature of 2 reduces accuracy at the end of training and increases loss as alpha is increased compared to a temperature of 1.
| Metrics | Control | Knowledge Distillation T = 1 | Knowledge Distillation T = 2 | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.420 ± 0.001 | 0.365 ± 0.001 | 0.260 ± 0.001 | 0.226 ± 0.000 | 0.310 ± 0.002 | 0.251 ± 0.001 | 0.221 ± 0.001 |
| Rank Disagreement | 0.997 ± 0.000 | 0.997 ± 0.000 | 0.997 ± 0.000 | 0.997 ± 0.000 | 0.997 ± 0.000 | 0.997 ± 0.000 | 0.996 ± 0.000 |
| Prediction Disagreement | 0.264 ± 0.003 | 0.256 ± 0.002 | 0.239 ± 0.002 | 0.235 ± 0.002 | 0.257 ± 0.002 | 0.258 ± 0.002 | 0.264 ± 0.002 |
| JS Divergence | 0.260 ± 0.001 | 0.221 ± 0.001 | 0.136 ± 0.001 | 0.106 ± 0.000 | 0.160 ± 0.001 | 0.101 ± 0.000 | 0.081 ± 0.000 |
| Accuracy | 0.680 ± 0.002 | 0.687 ± 0.002 | 0.700 ± 0.001 | 0.703 ± 0.002 | 0.685 ± 0.001 | 0.684 ± 0.002 | 0.678 ± 0.001 |
| Loss | 1.307 ± 0.009 | 1.342 ± 0.009 | 1.608 ± 0.015 | 1.833 ± 0.022 | 1.492 ± 0.014 | 1.725 ± 0.019 | 1.935 ± 0.019 |
| KD 0.1 | ✗ | ✓ | ✓ | ✗ | ✓ | ✗ |
| KD 0.5 | ✗ | ✓ | ✓ | ✗ | ✓ | ✗ |
| KD 0.9 | ✓ | ✓ | ✗ | ✓ | ✗ | ✗ |
Impact on Functional Similarity:
For a temperature of 2 we do see some deviation from temperature of 1 in regards to similarity, however not for prediction disagreement in Table 25 and 26). We observe that negative transfer (see Figure 13) is reduced but the negative asymmetry is preserved. Therefore, increasing the KD temperature reduces its overall utility when moving from a temperature of 1 to 2 (Table 25), while not eliminating the negative asymmetric transfer we uncover (Figure 13).
4.4.2 Language
For the below results, we use the Nano-GPT teacher with a Pico-GPT student, and study the nuances of temperature’s impact on knowledge transfer in the language domain.
| Metrics | Control | Knowledge Distillation T = 1 | Knowledge Distillation T = 2 | Knowledge Distillation T = 4 | ||||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.202 ± 0.000 | 0.198 ± 0.000 | 0.181 ± 0.000 | 0.172 ± 0.000 | 0.197 ± 0.000 | 0.183 ± 0.000 | 0.181 ± 0.000 | 0.199 ± 0.000 | 0.189 ± 0.000 | 0.193 ± 0.000 |
| Rank Disagreement | 0.915 ± 0.000 | 0.915 ± 0.000 | 0.912 ± 0.000 | 0.911 ± 0.000 | 0.907 ± 0.000 | 0.896 ± 0.000 | 0.892 ± 0.000 | 0.893 ± 0.000 | 0.880 ± 0.000 | 0.876 ± 0.000 |
| Prediction Disagreement | 0.252 ± 0.000 | 0.247 ± 0.000 | 0.226 ± 0.000 | 0.214 ± 0.000 | 0.250 ± 0.000 | 0.235 ± 0.000 | 0.230 ± 0.000 | 0.251 ± 0.001 | 0.244 ± 0.000 | 0.245 ± 0.000 |
| JS Divergence | 0.055 ± 0.000 | 0.054 ± 0.000 | 0.047 ± 0.000 | 0.043 ± 0.000 | 0.053 ± 0.000 | 0.047 ± 0.000 | 0.047 ± 0.000 | 0.054 ± 0.000 | 0.050 ± 0.000 | 0.051 ± 0.000 |
| Accuracy | 0.571 ± 0.000 | 0.572 ± 0.000 | 0.575 ± 0.000 | 0.574 ± 0.000 | 0.572 ± 0.000 | 0.572 ± 0.000 | 0.569 ± 0.000 | 0.570 ± 0.000 | 0.568 ± 0.000 | 0.562 ± 0.000 |
| Loss | 1.475 ± 0.001 | 1.471 ± 0.002 | 1.472 ± 0.001 | 1.496 ± 0.002 | 1.513 ± 0.003 | 1.571 ± 0.002 | 1.622 ± 0.002 | 1.528 ± 0.002 | 1.592 ± 0.002 | 1.663 ± 0.002 |
| KD 0.1 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
| KD 0.5 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
Impact on Accuracy:
In Table, 27, we only consider the results in the KD control with increasing temperature (1,2,4) against SIDDO. The raw results tables for these temperatures against RCD and SIDDO can be found in Appendix Section C.3.1. When considering accuracy and loss, increasing the temperature to 2 and 4 results in reduced accuracy increase when compared to using a temperature of 1, for all teacher seeds (0-2). For ease and clarity, the following analysis is provided for teacher seed 0; however, it holds for all teacher seeds. With teacher seed 0, the best accuracy (57.50%) was achieved with temperature 1, with an accuracy of 57.20% for temperature 2, and 57.00% for temperature 4 (see Table 27). It can be understood that the beneficial components of KD are reduced under the increased temperature; additionally, there is statistically supported but less functional knowledge passed to the student model when using a temperature of 2 and 4.
| KD 0.1 | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
| KD 0.5 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
Impact on Functional Similarity:
When increasing the teacher temperature from 1 to 4, the functional similarity distance between student and teacher models largely increases, resulting in less transfer at higher temperature values. Consistent with this reduction in knowledge transfer as temperature increases, we observe a reduction in maximum correct agreement (Figure 14): at temperature 1 it is 1.85%, at temperature 2 it is 0.85%, and at temperature 4 it is 0.11%. We also observe a reduction in maximum incorrect agreement: at temperature 1 it is 6.32%, at temperature 2 it is 3.80%, and at temperature 4 it is 2.20%. Figure 14 shows how the proportion of negative and positive transfer changes as temperature increases. Moreover, after adjusting for temperature, the negative asymmetric transfer we identify remains apparent and statistically supported regardless of temperature value.
This section demonstrates that using higher temperature values ultimately reduces the overall utility of KD’s knowledge transfer. As a result, the amount of knowledge transfer decreases, but when there is statistically supported transfer, we continue to observe negative asymmetric transfer. This follows logically as temperature : the outputs effectively converge to the uniform distribution (as in RCD), with all values equak to ( number of classes). This produces a constant output, which in turn negates the effect of distillation.
4.5 Adversarial Transfer Facilitated by Knowledge Distillation
This section shows that KD can match the teacher’s incorrect representations via a controlled adversarial experiment. We train an adversarial Nano-GPT teacher whose training set has every occurrence of ‘t’ ‘h’ ‘e’ replaced with ‘t’ ‘h’ ‘a’. Given the dataset’s Zipf’s law (Piantadosi, 2014) (Appendix Table 91, visualised in Figure 15), ‘e’ is the most likely character after ‘SPACE’. Therefore, if adversarial transfer is facilitated via knowledge transfer, a student trained with the adversarial teacher should predict ‘t’ ‘h’ ‘a’ more than ‘t’ ‘h’ ‘e’, compared to the controls trained without the adversarial teacher signal. Note that “tha” never naturally occurs in the dataset. To test whether incorrect knowledge can transfer asymmetrically from teacher to student, we use this setup to highlight potential safety concerns with knowledge distillation. In this case, the teacher has a known vulnerability and has been poisoned to predict an incorrect token. We show that this can be transferred to the student under standard distillation, resulting in a more substantial transfer of the teacher’s incorrect knowledge. If such a simple case of adversarial transfer can be engineered with minimal effort, then employing knowledge distillation in practice requires safety considerations. Since establishing when neural networks learn poor or harmful representations across scales remains an ongoing research endeavour, this experiment is crucial to understanding the impacts of negative asymmetric transfer in KD.
| Predicted Word | Teacher | SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 |
| tha | 454 | 105.900 4.168 | 106.000 3.046 | 199.100 13.391 | 436.200 7.984 | 104.600 3.898 | 114.800 3.056 | 126.900 8.068 |
| the | 285 | 665.100 7.675 | 675.500 10.228 | 583.400 17.536 | 343.600 6.358 | 668.800 12.713 | 712.500 12.480 | 826.300 20.203 |
On clean evaluation prompts containing “th_”, we measure how often models complete to “tha” vs. “the” and aggregate results per teacher seed, as seen for seed 0 in Table 30 (seeds 1-2 in Appendix Tables 93 and 94). KD, particularly at higher , markedly increases the rate of “tha” completions and suppresses “the” relative to both controls, demonstrating that KD can selectively copy a targeted error pattern even when overall behaviour appears benign. This experiment provides causal evidence that KD transmits specific erroneous structure, not merely broad functional alignment, sharpening the safety implications of our main findings: practitioners may unknowingly inherit unintended behaviours from the teacher, reinforcing our characterisation of KD as a data-dependent regulariser with a negative asymmetric payoff. Full details and per-seed statistics are provided in Appendix F.2.
5 Gradient-Level Explanation of Asymmetric Transfer
To summarise our findings, we provide a concise gradient-level mechanism that explains why distillation can preferentially transmit teacher errors when the teacher is imperfect. We emphasise that this section provides a mechanistic account consistent with the empirical negative asymmetric transfer observed throughout the paper, rather than claiming that such error transfer is a universally unavoidable phenomenon. In Appendix A we extend our functional analysis with information-theoretic and geometric perspectives to quantify when and how alignment with the teacher becomes harmful. These analyses connect the KD objective to a mechanism that can increase student agreement with teacher-incorrect predictions, and together with our targeted transfer experiment motivate auditing teacher error modes when distilling models used in safety-critical settings.
Consider the standard KD objective (with for clarity):
where and are the student and teacher logits, respectively. Using the standard softmax-cross-entropy identity and the KL gradient with respect to student logits, we obtain for each logit index :
with and .
When is the ground truth class (), the gradient includes both supervision and teacher alignment. When is an incorrect class (), the full gradient is:
In particular, the teacher-dependent component of the gradient on incorrect classes is (equivalently, the teacher-dependent term in the KD part is ). Thus, whenever the teacher assigns non-trivial probability mass to an incorrect class, the KD term exerts pressure for the student to allocate probability mass to that class. The strength of this teacher-driven component scales with and with the probability mass the teacher assigns to incorrect classes.
This derivation provides a mechanism consistent with our central finding: when the teacher is imperfect, KD disproportionately transfers errors to the student. The resulting alignment is asymmetric, favouring incorrect predictions; when the teacher matches training labels, with low loss, KD is negated.
By contrast, replacing the teacher with a class-uniform target – as in label smoothing (Appendix B) or our Random Control Distillation – then becomes constant across classes and removes teacher-specific structure from the KD term. In this case, the optimisation can still provide a smoothing/regularisation effect, but it no longer contains a teacher-shaped signal that differentially emphasises particular incorrect classes. Empirically, these controls often match or exceed KD in accuracy while not inducing the same pattern of increased agreement on teacher-incorrect behaviour.
Additionally, with temperature , the distillation term uses softened distributions and (with the usual prefactor in Eq. 1). Increasing reduces the sharpness of the teacher distribution, thereby reducing the concentration of probability mass on specific classes (including potentially incorrect ones), which can reduce measurable teacher–student functional transfer in practice. Consistent with our experiments in Section 4.4, increased typically attenuates transfer but does not necessarily eliminate the negative asymmetric payoff when transfer remains statistically supported. Overall, we argue that the observed asymmetric transfer in KD is not incidental: it is consistent with a mechanism that arises directly from the standard KD objective and can manifest across modalities, model sizes, and dataset scales whenever the teacher assigns non-trivial probability mass to incorrect classes. We empirically show that higher teacher loss exacerbates negative asymmetric transfer. Since regularisation methods that reduce training-set memorisation often also increase teacher loss, as observed with augmentation, our mechanistic account suggests that this phenomenon may persist across alternative training paradigms.
6 Safety Implications of Findings
Our results can be summarised into three key points: 1) knowledge distillation can enable statistically supported functional transfer; 2) the accuracy and loss benefits attributed to knowledge distillation are often matched or even exceeded by random controls; 3) knowledge distillation disproportionately transfers teacher-incorrect behaviour, with this asymmetry increasing as reliance on the teacher signal grows. Considering these findings – particularly points 2 and 3 – knowledge distillation raises practical safety concerns.
While it is often assumed that KD benefits student models, our results challenge this assumption by showing that undesirable teacher behaviour can be transferred to the student when transfer is statistically supported. We present a concrete case of targeted adversarial transfer facilitated by KD in Section 4.5. More broadly, our findings suggest that knowledge distillation is not uniformly reliable as a compression-by-transfer mechanism: at best it yields minimal positive functional transfer, and when transfer is substantial it amplifies teacher errors. This motivates auditing teacher error modes and evaluating distillation outcomes beyond accuracy/loss when KD is used in safety-relevant settings.
7 Conclusion
Across controlled self-distillation, small/large-scale settings, cross-modality (image, audio, language), a targeted error test, capacity scaling, standard KD with smaller students, and multiple KD variants, KD seldom delivers robust “knowledge transfer” as measured by teacher–student functional similarity beyond counterfactual controls. When transfer occurs, it is often marginal and inconsistent, and increases with teacher imperfection, amplifying the teacher’s errors more than its correct behaviour (negative asymmetry). By contrast, RCD often yields the best loss/accuracy, indicating that reported gains can arise from generic regularisation rather than faithful knowledge transmission. Our targeted language experiment confirms that KD can copy specific erroneous patterns, and our scaling law experiments show that increasing student capacity amplifies incorrect agreement faster than correct. Overall, our functional analysis links these empirical patterns to the structure of the standard distillation gradient (Section 5), providing a mechanistic account of why teacher error mass can drive negative asymmetric transfer under logit matching. We therefore reframe KD as a data-dependent regulariser whose transfer component can be negative and asymmetric, with clear safety implications: audit teacher error structure and report functional transfer analyses (teacher-correct vs. teacher-incorrect agreement) alongside accuracy/loss.
8 Limitations
Our conclusions are drawn from a large but finite set of datasets, architectures, and training recipes, and may not capture regimes with substantially different optimisation. Our statistical definition of transfer is functional (teacher–student agreement beyond counterfactual controls) and therefore does not measure all possible forms of “knowledge” (e.g., internal feature reuse). The gradient-level explanation (Section 5) focuses on standard logit-matching KD; although we observe negative asymmetric transfer under feature-map matching, we do not provide a corresponding mechanistic derivation for that variant. Extending this analysis to broader distillation objectives is a natural direction for future work.
Appendix A Extended Functional Analysis: Information-Theoretic and Geometric Perspectives
We apply two additional metrics: Variation of Information (VoI), an information-theoretic measure over discrete labellings that penalises confident mispredictions (Meilă, 2003), and Orthogonal Procrustes Distance (OPD), a geometric alignment metric over output representations (Schönemann, 1966; Ding et al., 2021). We compute VoI and OPD for two representative setups: ResNet18 on SVHN and ResNet50 on TinyImageNet (teacher seed 0). OPD closely tracks trends observed in Activation Distance and JS Divergence, showing decreasing student–teacher discrepancy as increases. VoI generally follows this trend, but diverges in specific cases (high on SVHN) where it increases despite stronger overall functional alignment. This is not contradictory: VoI is sensitive to changes in the discrete predicted labels, and therefore increases when alignment is driven by agreement on teacher-incorrect predictions. Its rise coincides with the strongest observed increase in student–teacher agreement on incorrect predictions, providing further evidence for KD’s negative asymmetric payoff. Overall, OPD confirms alignment, while VoI highlights when that alignment corresponds to the transfer of incorrect information. Moreover, this behaviour is consistent with our gradient-based analysis: the per-logit gradient under KD pulls the student toward the teacher’s incorrect predictions with strength proportional to and to the teacher’s own loss. VoI captures the cost of absorbing these errors, providing an explicit signal of negative information transfer. OPD, meanwhile, captures alignment that may or may not be beneficial.
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.063 0.002 | 0.064 0.001 | 0.060 0.001 | 0.059 0.001 | 0.144 0.001 | 0.493 0.000 | 0.849 0.000 |
| Rank Disagreement | 0.696 0.003 | 0.688 0.004 | 0.684 0.003 | 0.681 0.003 | 0.800 0.002 | 0.798 0.002 | 0.802 0.003 |
| Prediction Disagreement | 0.045 0.001 | 0.046 0.001 | 0.043 0.001 | 0.042 0.001 | 0.042 0.001 | 0.043 0.001 | 0.046 0.001 |
| JS Divergence | 0.025 0.001 | 0.025 0.001 | 0.023 0.001 | 0.022 0.000 | 0.053 0.000 | 0.201 0.000 | 0.431 0.000 |
| Information Variation | 0.550 0.051 | 0.588 0.049 | 0.594 0.024 | 0.614 0.018 | 0.638 0.000 | 0.638 0.000 | 0.638 0.000 |
| Procrustes Distance | 0.165 0.003 | 0.168 0.004 | 0.164 0.003 | 0.162 0.005 | 0.291 0.001 | 0.304 0.001 | 0.311 0.003 |
| Accuracy | 0.952 0.001 | 0.951 0.001 | 0.954 0.001 | 0.954 0.001 | 0.957 0.001 | 0.957 0.001 | 0.955 0.001 |
| Loss | 0.385 0.011 | 0.344 0.008 | 0.310 0.006 | 0.293 0.004 | 0.236 0.003 | 0.692 0.001 | 1.698 0.001 |
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Information Variation | 0.519 0.017 | 0.520 0.017 | 0.518 0.022 | 0.533 0.014 | 0.856 0.002 | 0.897 0.001 | 0.907 0.002 |
| Procrustes Distance | 0.050 0.000 | 0.050 0.000 | 0.050 0.000 | 0.049 0.000 | 0.433 0.000 | 0.664 0.000 | 0.553 0.000 |
Appendix B Random Control Distillation (RCD) Comparison to Label Smoothing
@@
We also rely on RCD, which retains the distillation form but replaces the teacher distribution with a class-uniform target ().
RCD therefore preserves the label-smoothing effect while removing teacher-specific semantic content. Across all metrics, we find that LS and RCD match or exceed KD in accuracy, yet exhibit no increase in functional similarity with the teacher, particularly on incorrect predictions when accuracy using RCD is preserved. This supports the interpretation that KD’s asymmetric error transfer is driven by the specific structure of the teacher logits, rather than softening per se.
@@
Appendix C Knowledge Distillation to Smaller Student
In this section we present the raw results for the other teacher seeds explored in this paper (1-2) to show that our findings are not the effect of a single seed but are general across teacher models and training seeds.
C.1 TinyImageNet ResNet50 Teacher to ResNet18 Student
@@
C.2 ImageNet ResNet50 Teacher to ResNet18 Student
@@
A pre-trained ResNet50 model taken from PyTorch with a top-1-accuracy of 80.858 and a top-5-accuracy of 95.434222https://docs.pytorch.org/vision/main/models/generated/torchvision.models.resnet50.html#torchvision.models.ResNet50_Weights is used as the teacher. As Pytorch only provides one set of pre-trained model weights there is only one teacher seed for this experiment. The ResNet18 student was trained on ImageNet (Deng et al., 2009) using the FFCV setup (Leclerc et al., 2023), where 100% of the training images were compressed to a JPEG with 90% quality. The data was normalized with a mean of (0.485, 0.456, 0.406) and a standard deviation of (0.229, 0.224, 0.225). The model utilised BlurPools (Zhang, 2019) within the convolutional layers, and was trained for 56 epochs, with a batch size of 1024 using SGD, momentum of 0.9, weight decay of 5e-5, a learning rate of 0.5 using a cyclic scheduler with a learning rate step ratio of 0,1 and step length of 30. The learning rate peak was at epoch 2. The input resolution started at 160 by 160, and started to ramp up to 192 by 192 at epoch 41 and ended at 192 by 192 at epoch 48.
C.2.1 The Effect Of Temperature
This section presents the raw results for temperature of 2 on the ImageNet dataset on seed 0, due to computational constraints this test was only conducted on seed 0.
| SIDDO | KD 0.1 | KD 0.5 | KD 0.9 | Rand KD 0.1 | Rand KD 0.5 | Rand KD 0.9 | |
| Activation Distance | 0.420 0.001 | 0.310 0.001 | 0.251 0.001 | 0.221 0.001 | 0.304 0.001 | 0.247 0.001 | 0.282 0.001 |
| Rank Disagreement | 0.997 0.000 | 0.997 0.000 | 0.997 0.000 | 0.996 0.000 | 0.997 0.000 | 0.997 0.000 | 0.997 0.000 |
| Prediction Disagreement | 0.263 0.002 | 0.257 0.001 | 0.256 0.001 | 0.264 0.001 | 0.259 0.001 | 0.273 0.001 | 0.310 0.001 |
| JS Divergence | 0.259 0.000 | 0.160 0.000 | 0.101 0.000 | 0.081 0.000 | 0.151 0.001 | 0.096 0.000 | 0.116 0.001 |
| Accuracy | 0.681 0.001 | 0.685 0.001 | 0.685 0.001 | 0.679 0.001 | 0.684 0.001 | 0.671 0.001 | 0.641 0.001 |
| Loss | 1.303 0.006 | 1.492 0.008 | 1.718 0.012 | 1.922 0.013 | 1.532 0.009 | 1.919 0.015 | 3.022 0.014 |
C.3 TinyShakespeare Nano-GPT Teacher to Pico-GPT Student